OpenStudy (anonymous):

Mike is building a rectangular pen for his dog. Which measurement does Mike need to find so that he knows how much fence material to buy? area perimeter volume circumference

OpenStudy (anonymous):

A rectangle has a length L and a width W. Area = L * W; Perimeter = 2 * (L + W) do you have numbers?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Try Area. usually area is used for a triangle i think. but try that if thats not correct go with perimeter, the last 2 you can cancel out because those are gonna be incorrect.